{2x+y=3 and 4x+2y=5}
Answer by Alan3354(69443)

2x+y=3
4x+2y = 5
4x+2y = 6 1st eqn time 2
-------------------------------- Subtract
0 = -1
--> inconsistent
No solution
If you graph them, you get 2 parallel lines, no intersection.